Copper and aluminum are being considered for a high-voltagetransmission line that must carry a current of 50.0 A. The resistance per unit length is to be0.200 /km. The densities of copper and aluminum are 8960 and 2700kg/m3, respectively.
Compute:
(a) the magnitude J of the averagecurrent density for a copper cable
1 A/m2

(b) the mass per unit length λ for a coppercable.
2 kg/m

(c) the magnitude J of the average current density for analuminum cable
3 A/m2

(d) the mass per unit length λ for an alumniumcable.
4 kg/m
Answer #1
Concepts and reason

The required concepts to solve these questions are current density, resistivity of conductor and electric field of conductor.

In part (a) calculate the current density for copper cable with the help of electric field equation and then in part (b) calculate the mass per unit length for the copper cable by using resistivity equation. In part (c), calculate the current density for aluminum cable and in part (d), calculate the mass per unit length for the aluminum cable by using mass density of aluminum equation.

Fundamentals

The expression for current density in term of electric field is,

J=EρJ = \frac{E}{\rho }

Here, EE is the electric field and ρ\rho is the resistivity of the conductor.

The expression for electric field of conductor is,

E=VLE = \frac{V}{L}

Here, VV is the potential difference and LL is the length of the conductor.

The expression for resistivity of the conductor is,

ρ=RAL\rho = \frac{{RA}}{L}

Here, AA is the area and RR is the resistance of the conductor.

The general equation of voltage is,

V=iRV = iR

Here, ii is the current and RR is the resistance.

(a)

The expression for electric field of conductor is,

E=VLE = \frac{V}{L}

The general equation of voltage is,

V=iRV = iR

The expression for current density in term of electric field is,

J=EρJ = \frac{E}{\rho }

Substitute (VL)\left( {\frac{V}{L}} \right) for EE in the above equation of current density.

J=VρLJ = \frac{V}{{\rho L}} …… (1)

Substitute iRiR for VV in equation (1).

J=iRρLJ = \frac{{iR}}{{\rho L}} …… (2)

Substitute 1.69×108Ωm1.69 \times {10^{ - 8}}{\rm{ }}\Omega \cdot {\rm{m}} for ρ\rho , 50A50{\rm{ A}} for ii and 0.200Ωkm10.200{\rm{ }}\Omega \cdot {\rm{k}}{{\rm{m}}^{ - 1}} for (RL)\left( {\frac{R}{L}} \right) in the equation (2).

J=(50A)(0.200Ωkm(1km103m))(1.69×108Ωm)=5.91×105Am2\begin{array}{c}\\J = \frac{{\left( {50{\rm{ A}}} \right)\left( {0.200{\rm{ }}\frac{\Omega }{{{\rm{km}}}}\left( {\frac{{1{\rm{ km}}}}{{{{10}^3}{\rm{ m}}}}} \right)} \right)}}{{\left( {1.69 \times {{10}^{ - 8}}{\rm{ }}\Omega \cdot {\rm{m}}} \right)}}\\\\ = 5.91 \times {10^5}{\rm{ A}} \cdot {{\rm{m}}^{ - 2}}\\\end{array}

(b)

The expression for resistivity of the conductor is,

ρ=RAL\rho = \frac{{RA}}{L}

Rearrange the equation in term of AA .

A=ρLRA = \frac{{\rho L}}{R}

The expression for mass density of copper is,

m=dVm = dV

Here, VV is the density of the copper.

Substitute ALAL for VV in equation of the mass density of copper.

m=d(AL)mL=dA\begin{array}{l}\\m = d\left( {AL} \right)\\\\\frac{m}{L} = dA\\\end{array}

Here, λ\lambda is used for (mL)\left( {\frac{m}{L}} \right) mass per unit length.

Substitute ρLR\frac{{\rho L}}{R} for AA and λ\lambda for mL\frac{m}{L} in above equation.

λ=dρ(RL)\lambda = d\frac{\rho }{{\left( {\frac{R}{L}} \right)}} …… (3)

Substitute 0.200Ωkm10.200{\rm{ }}\Omega \cdot {\rm{k}}{{\rm{m}}^{ - 1}} for (RL)\left( {\frac{R}{L}} \right) , 8960kgm38960{\rm{ kg}} \cdot {{\rm{m}}^{ - 3}} for dd and 1.69×108Ωm1.69 \times {10^{ - 8}}{\rm{ }}\Omega \cdot {\rm{m}} in above equation.

λ=(8960kgm3)(1.69×108Ωm0.200Ωkm(1km103m))=(8960kgm3)(1.69×108Ωm0.200×103Ωm1)=0.757kgm1\begin{array}{c}\\\lambda = \left( {8960{\rm{ kg}} \cdot {{\rm{m}}^{ - 3}}} \right)\left( {\frac{{1.69 \times {{10}^{ - 8}}{\rm{ }}\Omega \cdot {\rm{m}}}}{{0.200{\rm{ }}\frac{\Omega }{{{\rm{km}}}}\left( {\frac{{1{\rm{ km}}}}{{{{10}^3}{\rm{ m}}}}} \right)}}} \right)\\\\ = \left( {8960{\rm{ kg}} \cdot {{\rm{m}}^{ - 3}}} \right)\left( {\frac{{1.69 \times {{10}^{ - 8}}{\rm{ }}\Omega \cdot {\rm{m}}}}{{0.200 \times {{10}^{ - 3}}{\rm{ }}\Omega \cdot {{\rm{m}}^{ - 1}}}}} \right)\\\\ = 0.757{\rm{ kg}} \cdot {{\rm{m}}^{ - 1}}\\\end{array}

(c)

Using the equation (2) current density for aluminum cable is,

J=iRρLJ = \frac{{iR}}{{\rho L}}

Here, ρ\rho is the resistivity of the aluminum cable.

Substitute 2.82×108Ωm2.82 \times {10^{ - 8}}{\rm{ }}\Omega \cdot {\rm{m}} for ρ\rho , 50A50{\rm{ A}} for ii and 0.200Ωkm10.200{\rm{ }}\Omega \cdot {\rm{k}}{{\rm{m}}^{ - 1}} for (RL)\left( {\frac{R}{L}} \right) in the above equation.

J=(50A)(0.200Ωkm(1km103m))(2.82×108Ωm)=3.5×105Am2\begin{array}{c}\\J = \frac{{\left( {50{\rm{ A}}} \right)\left( {0.200{\rm{ }}\frac{\Omega }{{{\rm{km}}}}\left( {\frac{{1{\rm{ km}}}}{{{{10}^3}{\rm{ m}}}}} \right)} \right)}}{{\left( {2.82 \times {{10}^{ - 8}}{\rm{ }}\Omega \cdot {\rm{m}}} \right)}}\\\\ = 3.5 \times {10^5}{\rm{ A}} \cdot {{\rm{m}}^{ - 2}}\\\end{array}

(d)

Using the equation (3) mass per unit length for aluminum cable is,

λ=dρ(RL)\lambda = d\frac{\rho }{{\left( {\frac{R}{L}} \right)}}

Here, ρ\rho is the resistivity and is the density of the aluminum cable.

Substitute 0.200Ωkm10.200{\rm{ }}\Omega \cdot {\rm{k}}{{\rm{m}}^{ - 1}} for (RL)\left( {\frac{R}{L}} \right) , 2700kgm32700{\rm{ kg}} \cdot {{\rm{m}}^{ - 3}} for dd and 2.82×108Ωm2.82 \times {10^{ - 8}}{\rm{ }}\Omega \cdot {\rm{m}} in above equation.

λ=(2700kgm3)(2.82×108Ωm0.200Ωkm(1km103m))=(2700kgm3)(2.82×108Ωm0.200×103Ωm1)=0.380kgm1\begin{array}{c}\\\lambda = \left( {2700{\rm{ kg}} \cdot {{\rm{m}}^{ - 3}}} \right)\left( {\frac{{2.82 \times {{10}^{ - 8}}{\rm{ }}\Omega \cdot {\rm{m}}}}{{0.200{\rm{ }}\frac{\Omega }{{{\rm{km}}}}\left( {\frac{{1{\rm{ km}}}}{{{{10}^3}{\rm{ m}}}}} \right)}}} \right)\\\\ = \left( {2700{\rm{ kg}} \cdot {{\rm{m}}^{ - 3}}} \right)\left( {\frac{{2.82 \times {{10}^{ - 8}}{\rm{ }}\Omega \cdot {\rm{m}}}}{{0.200 \times {{10}^{ - 3}}{\rm{ }}\Omega \cdot {{\rm{m}}^{ - 1}}}}} \right)\\\\ = 0.380{\rm{ kg}} \cdot {{\rm{m}}^{ - 1}}\\\end{array}

Ans: Part a

The magnitude JJ of the current density for a copper cable is 5.91×105Am25.91 \times {10^5}{\rm{ A}} \cdot {{\rm{m}}^{ - 2}} .

Part b

The mass per unit length λ\lambda for a copper cable is 0.757kgm10.757{\rm{ kg}} \cdot {{\rm{m}}^{ - 1}} .

Part c

The magnitude JJ of the current density for an aluminum cable is 3.5×105Am23.5 \times {10^5}{\rm{ A}} \cdot {{\rm{m}}^{ - 2}} .

Part d

The mass per unit length λ\lambda for an aluminum cable is 0.380kgm10.380{\rm{ kg}} \cdot {{\rm{m}}^{ - 1}} .